A glass of relative permitivity 4 is kept in an external electric field of magnitude 10^2 Vm^-1. Calculate the polarization vector, molecular/ atomic polarisability and the reflective index of the glass.
1
Expert's answer
2020-10-20T07:06:08-0400

Polarization vector, "P=\chi\epsilon_{o}E"

"\epsilon=1+\chi"

"\chi=\epsilon-1=4-1=3"

Therefore, "P=\chi\epsilon_{o}E=3\times8.85\times10^{-12}\times100=26.55\times10^{-10}" "\frac{FV}{m^{2}}"


Molecular/atomic polarisability, "P" "=N\alpha\epsilon_{o}E"

"\alpha=\frac{P}{N\epsilon_{o}E}= \frac{26.55\times10^{-10}}{6.02\times10^{23}\times8.85\times10^{-12}\times100}" ="5.0\times10^{-24}m^{3}"


the refractive index, "\eta" "=\sqrt\epsilon\mu = \sqrt(4\times0.99)=1.99"